# On circumcentered direct methods for monotone variational inequality problems

### v1

@@ 4 An explicit circumcenter method for monotone variational inequality problem
$g(x)=\max_{1\leq i\leq m}g_{i}(x).$ (4.5)
Set
$\tilde{y}^{k}:=y^{k,j(k)}.$ (4.6)
Define $\eta_{k}:=\max\{1,\|T(\tilde{y}^{k})\|\}$. Take
$z^{k+1}:=T_{\mathbf{S}_{k},\mathbf{D}}\left(\tilde{y}^{k}-\frac{\beta_{k}}{\eta_{k}}T(\tilde{y}^{k})\right).$ (4.7)
If $z^{k+1}=y^{k}$, stop. Otherwise, define
$\sigma_{k}:=\sum_{j=0}^{k}\frac{\beta_{k}}{\eta_{k}}=\sigma_{k-1}+\frac{\beta_{k}}{\eta_{k}},$ (4.8)
$x^{k+1}:=\left(1-\frac{\beta_{k}}{\eta_{k}\sigma_{k}}\right)x^{k}+\frac{\beta_{k}}{\eta_{k}\sigma_{k}}\tilde{y}^{k}.$ (4.9)
We assume in this section that there exist a Slater point $w\in\mathds{R}^{n}$. This assumption is only needed in the inner loop (4.3) in order to check whether the current iterate in the loop is close enough to the feasible set.
4.1 Analysis of the direct method
For the convergence of our method, we assume that $T$ is monotone and continuous. Observe that $\partial g_{i}(x)\neq\emptyset$ for $x\in\mathds{R}^{n}$ for all $i\in\{1,\ldots,m\}$, because we assume that $g_{i}$ is convex and dom$(g_{i})=\mathds{R}^{n}$, for all $i\in\{1,\ldots,m\}$. First, we establish that Algorithm A is well-defined.
@@ 4.1 Analysis of the direct method
  • ii)
    If $z^{k+1}=\tilde{y}^{k}$, then $\tilde{y}^{k}\in{\rm SOL}(F,C)$.
  • iii)
    $j(k)$ is well defined.
@@
  • i)
    By the definition of $y^{k,j+1}$ and Proposition 2.3, we know that $y^{k+1}\in C_{i}$ for all $i\in\{1,\ldots,m\}$. According to the definition of $C_{i}$ and $g$, we know that (i) holds automatically.
  • ii)
    Suppose that $z^{k+1}=\tilde{y}^{k}$. Then, by Proposition 2.3, we get that $\mathbf{\tilde{y}}^{k}\in\mathbf{H}^{k}\subset\mathbf{S}^{k}\subset\mathds{R}^{nm}$. Therefore, $\tilde{y}^{k}\in S_{i}^{k}$ for all $i\in\{1,\ldots,m\}$. Given $i\in\{1,\ldots,m\}$, we have two cases, according to the definition of $S_{i}^{k}$ in (2.4). If $S_{i}^{k}=C_{i}$, then we have $z^{k+1}=\tilde{y}^{k}\in C_{j}$. Otherwise, we have $g_{j}(\tilde{y}^{k})\leq g_{j}(\tilde{y}^{k})+\langle v^{jk},z^{k+1}-\tilde{y}^{k}\rangle\leq 0$, i.e., $\tilde{y}^{k}\in C_{j}$. By the above argument above, we conclude that for any $i\in\{1,\ldots,m\}$, we have $\tilde{y}^{k}\in C_{i}$.
  • iii)
    Assume by contradiction that $\frac{g(y^{k,j})\|y^{k,j}-w\|}{g(y^{k,j})-g(w)}>\theta\beta_{k}$ for all $j$. Thus, we obtain an infinite sequence $\{y^{k,j}\}$ such that
    $\liminf_{j\rightarrow\infty}\frac{g(y^{k,j})\|y^{k,j}-w\|}{g(y^{k,j})-g(w)}\geq\theta\beta_{k}>0.$ (4.10)
    Taking into account the inner loop in $j$ given in (4.3) i.e., $y^{k,j+1}:=T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j})$ for each fixed $k$, we obtain, for each $x\in C$,
    $\begin{split}\|y^{k,j+1}-x\|^{2}&=\|T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j})-T_{\mathbf{S}_{k},\mathbf{D}}(x)\|^{2}=\|P_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j})-P_{\mathbf{S}_{k},\mathbf{D}}(x)\|\\ &\leq\|y^{k,j}-x\|^{2}-\|y^{k,j+1}-y^{k,j}\|^{2}\\ &\leq\|y^{k,j}-x\|^{2},\end{split}$ (4.11)
    where the first equality follows from Proposition 2.3 and the first inequality holds because of the firm nonexpansiveness of the projection operator. Thus, $\{y^{k,j}\}_{j\geq 0}$ is a quasi-Fejér convergent to $C$, and hence it is bounded by Proposition 2.15. It follows that
    $\tau:=\frac{1}{-g(w)}\sup_{0\leq j\leq\infty}\|y^{k,j}-w\|,$ (4.12)
@@
    $\lim_{j\rightarrow\infty}\|y^{k,j+1}-y^{k,j}\|=\lim_{j\rightarrow\infty}\|T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j})-y^{k,j}\|=0.$ (4.14)
@@
    $\liminf_{j\rightarrow\infty}g(y^{k,j})\leq 0.$ (4.16)
@@
    $\begin{split}\liminf_{j\rightarrow\infty}\frac{g(y^{k,j})\|y^{k,j}-w\|}{g(y^{k,j})-g(w)}&\leq\liminf_{j\rightarrow\infty}\frac{g(y^{k,j})\|y^{k,j}-w\|}{-g(w)}\\ &\leq\frac{1}{-g(w)}\sup_{0\leq j\leq\infty}\|y^{k,j}-w\|\liminf_{j\rightarrow\infty}g(y^{k,j})\\ &=\tau\liminf_{j\rightarrow\infty}g(y^{k,j})\leq 0,\end{split}$ (4.17)
    contradicting (4.10). Therefore, $j(k)$ is well defined.
@@
If ${\rm SOL}(F,C)$ is nonempty, then $\{\tilde{y}^{k}\}$ and $\{z^{k}\}$ are quasi-Fejer convergent to ${\rm SOL}(F,C)$.
@@
$\frac{\|F(\tilde{y}^{k})\|}{\eta_{k}}\leq 1.$ (4.19)
Take $x^{*}\in{\rm SOL}(F,C)$. Then,
$\begin{split}\|\tilde{y}^{k}-x^{*}\|^{2}&=\|y^{k,j(k)}-x^{*}\|^{2}=\|T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j(k-1)})-T_{\mathbf{S}_{k},\mathbf{D}}(x^{*})\|^{2}\leq\|y^{k,j(k)-1}-x^{*}\|^{2}\\ &=\|T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j(k)-2})-T_{\mathbf{S}_{k},\mathbf{D}}(x^{*})\|^{2}\leq\|y^{k,j(k)-2}-x^{*}\|^{2}\\ &\cdots\\ &\leq\|y^{k,1}-x^{*}\|^{2}=\|T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,0})-T_{\mathbf{S}_{k},\mathbf{D}}(x^{*})\|^{2}\\ &\leq\|y^{k,0}-x^{*}\|^{2}=\|z^{k}-x^{*}\|^{2}.\end{split}$ (4.20)
@@
$\begin{split}\|\tilde{y}^{k+1}-x^{*}\|^{2}&\leq\|z^{k+1}-x^{*}\|^{2}=\left\|T_{\mathbf{S}_{k},\mathbf{D}}(\tilde{y}^{k}-\frac{\beta_{k}}{\eta_{k}}F(\tilde{y}^{k}))-T_{\mathbf{S}_{k},\mathbf{D}}(x^{*})\right\|^{2}\leq\|\tilde{y}^{k}-\frac{\beta_{k}}{\eta_{k}}F(\tilde{y}^{k})-x^{*}\|^{2}\\ &=\|\tilde{y}^{k}-x^{*}\|^{2}+\frac{\|F(\tilde{y}^{k})\|^{2}}{\eta_{k}^{2}}\beta_{k}^{2}-2\frac{\beta_{k}}{\eta_{k}}\langle F(\tilde{y}^{k}),\tilde{y}^{k}-x^{*}\rangle\\ &\leq\|\tilde{y}^{k}-x^{*}\|^{2}+\beta_{k}^{2}-2\frac{\beta_{k}}{\eta_{k}}\langle F(x^{*}),\tilde{y}^{k}-x^{*}\rangle\\ &=\|\tilde{y}^{k}-x^{*}\|^{2}+\beta_{k}^{2}-2\frac{\beta_{k}}{\eta_{k}}(\langle F(x^{*}),\tilde{y}^{k}-P_{C}(\tilde{y}^{k})\rangle+\langle F(x^{*}),P_{C}(\tilde{y}^{k})-x^{*}\rangle)\\ &\leq\|\tilde{y}^{k}-x^{*}\|^{2}+\beta_{k}^{2}-2\frac{\beta_{k}}{\eta_{k}}\|F(x^{*})\|\|P_{C}(\tilde{y}^{k})-\tilde{y}^{k}\|\\ &\leq|\tilde{y}^{k}-x^{*}\|^{2}+\tilde{\theta}\beta_{k}^{2}\\ &\leq\|z^{k}-x^{*}\|^{2},\end{split}$ (4.21)
using (4.20) in the last inequality, the nonexpansiveness of $T_{\mathbf{S}_{k},\mathbf{D}}$ in the second one, the monotonicity of $F$ and (4.19) in the third one, $x^{*}\in{\rm SOL}(F,C)$ in the fourth one, Cauchy-Schwartz inequality in the fifth one, Lemma 2.18 and definition of $j(k)$ in the sixth one, and (4.20) in the last one.
In view of Definition 2.12, we conclude that $\{\tilde{y}^{k}\}$ and $\{z^{k}\}$ are quasi-Fejér convergent to ${\rm SOL}(F,C)$. ∎
@@
Let $\{z^{k}\}$, $\{\tilde{y}^{k}\}$ and $\{x^{k}\}$ be the sequences generated by Algorithm 2. Assume that ${\rm SOL}(F,C)$ is nonempty. Then
  • i)
@@
  • iii)
    $\lim_{k\rightarrow\infty}dist(x^{k},C)=0$;
  • iv)
@@
    $\begin{split}\sigma_{k}x^{k+1}=\left(\sigma_{k}-\frac{\beta_{k}}{\eta_{k}}\right)x^{k}+\frac{\beta_{k}}{\eta_{k}}\tilde{y}^{k}&=\sigma_{k-1}x^{k}+\frac{\beta_{k}}{\eta_{k}}\tilde{y}^{k}\\ &=\sigma_{k-2}x^{k-1}+\frac{\beta_{k-1}}{\eta_{k-1}}\tilde{y}^{k-1}+\frac{\beta_{k}}{\eta_{k}}\tilde{y}^{k}\\ &\cdots\\ &=\sigma_{0}x^{1}+\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}\tilde{y}^{j}\\ &=\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}\tilde{y}^{j},\end{split}$ (4.22)
@@
  • ii)
    For $\{\tilde{y}^{k}\}$ we use Lemma 4.2 and Proposition 2.15(i). For $\{F(\tilde{y}^{k})\}$, we use the boundedness of $\{\tilde{y}^{k}\}$ and the fact that $F$ is monotone and continuous. For $\{x^{k}\}$, use the boundedness of $\{\tilde{y}^{k}\}$ and (i).
  • iii)
@@
    $dist(\tilde{y}^{k},C)\leq\theta\beta_{k}.$ (4.23)
@@
    $\tilde{x}^{k+1}:=\frac{1}{\sigma_{k}}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}P_{C}(\tilde{y}^{k}).$ (4.24)
@@
    $\tilde{\beta}:=\sum_{j=0}^{\infty}\beta_{j}^{2}.$ (4.25)
@@
    $\begin{split}dist(x^{k+1},C)&\leq\|x^{k+1}-\tilde{x}^{k+1}\|=\left\|\frac{1}{\sigma_{k}}\left(\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}(\tilde{y}^{j}-P_{C}(\tilde{y}^{j}))\right)\right\|\\ &\leq\frac{1}{\sigma_{k}}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}\|\tilde{y}^{j}-P_{C}(\tilde{y}^{j})\|\\ &\leq\frac{\theta}{\sigma_{k}}\sum_{j=0}^{k}\frac{\beta_{j}^{2}}{\eta_{j}}\leq\frac{\theta}{\sigma_{k}}\sum_{j=0}^{k}\beta_{j}^{2}\leq\theta\frac{\tilde{\beta}}{\sigma_{k}},\end{split}$ (4.26)
[…]
@@
    $\lim_{k\rightarrow\infty}\sigma_{k}=\lim_{k\rightarrow\infty}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}\geq\lim_{k\rightarrow\infty}\frac{1}{\gamma}\sum_{j=0}^{k}\beta_{j}=\infty,$ (4.27)
    using the fact that $\eta_{j}=max\{1,\|F(\tilde{y})\|\}\leq max\{1,\gamma\}\leq\gamma$ for all $j$ in the inequality, and (4.1) in the first equality. Thus, taking limits in (4.26), we get
    $\lim_{k\rightarrow\infty}dist(x^{k},C)=0,$
    establishing (iii). Item (iv) follows immediately from (iii).
Theorem 4.4.
If ${\rm SOL}(F,C)$ is nonempty then all the cluster points of $\{x^{k}\}$ generated by Algorithm 2 are solutions of ${\rm VIP}(F,C)$.
@@
$\begin{split}\|z^{j+1}-x\|&=\left\|T_{\mathbf{S}_{k},\mathbf{D}}\left(\tilde{y}^{j}-\frac{\beta_{j}}{\eta_{j}}F(\tilde{y}^{j})\right)-T_{\mathbf{S}_{k},\mathbf{D}}(x)\right\|\leq\left\|\left(\tilde{y}^{j}-\frac{\beta_{j}}{\eta_{j}}F(\tilde{y}^{j})\right)-x\right\|\\ &=\|\tilde{y}^{j}-x\|^{2}+\frac{\|F(\tilde{y}^{j})\|}{\eta_{j}^{2}}\beta_{j}^{2}-2\frac{\beta_{j}}{\eta_{j}}\langle F(\tilde{y}^{j}),\tilde{y}^{j}-x\rangle\\ &\leq\|\tilde{y}^{j}-x\|^{2}+\beta_{j}^{2}+2\frac{\beta_{j}}{\eta_{j}}\langle F(x),x-\tilde{y}^{j}\rangle,\end{split}$ (4.28)
using nonexpansiveness of $T_{\mathbf{S}_{k},\mathbf{D}}$ in the first inequality, and monotonicity of $F$ and (4.19) in the last inequality. Summing (4.28) from 00 to $k-1$ and dividing by $\sigma_{k-1}$, we obtain, in view of Lemma 4.3(ii),
$\frac{(\|z^{k}-x\|^{2}-\|z^{0}-x\|^{2})}{\sigma_{k-1}}\leq\frac{\sum_{j=0}^{k-1}\beta_{j}^{2}}{\sigma_{k-1}}+\langle F(x),x-x^{k}\rangle.$ (4.29)
@@
for all $x\in C$. Using Proposition 2.14, we get that $\hat{x}\in{\rm SOL}(F,C)$. Therefore, all the cluster points of $\{x^{k}\}$ solve ${\rm VIP}(F,C)$. ∎

### v2

@@ 4. An explicit circumcenter method for the monotone variational inequality problem
(4.3) $g(x)=\max_{1\leq i\leq m}g_{i}(x).$
     if Condition is not met then $j\leftarrow j+1$
     end if
   until Condition (4.2) is met
   Set $j(k)\leftarrow j$ and define:
(4.4) $\tilde{y}^{k}:=y^{k,j(k)}.$
  end if
  Define $\eta_{k}\coloneqq\max\{1,\|F(\tilde{y}^{k})\|\}$.
  Compute next $z$ iterate:
(4.5) $z^{k+1}:=T_{\mathbf{S}_{k},\mathbf{D}}\left(\tilde{y}^{k}-\frac{\beta_{k}}{\eta_{k}}F(\tilde{y}^{k})\right).$
  if $z^{k+1}=\tilde{y}^{k}$ then
   stop
  end if
  Update partial sum and average iterate:
(4.6) $\sigma_{k}:=\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}.\quad(\text{accumulated})$
(4.7) $x^{k+1}:=\left(1-\frac{\beta_{k}}{\eta_{k}\sigma_{k}}\right)x^{k}+\frac{\beta_{k}}{\eta_{k}\sigma_{k}}\tilde{y}^{k}.$
end for
We assume that a Slater point $w\in\mathds{R}^{n}$ is available. This assumption is needed only in the inner loop (4.1) in order to certify whether the current iterate in the loop is sufficiently close to the feasible set.
Remark 4.1.
The computation in the inner loop consists solely of circumcenter steps, which require projections onto halfspaces and are therefore quite inexpensive. Even if the evaluation of $F$ is costly, it is performed only once per iteration, as in the algorithms proposed in [49].
4.1. Analysis of the direct method
In this subsection we assume that $F,C$ satisfy Assumptions 1.5 and 1.10, i.e. $\operatorname{VIP}(F,C)$ has solutions and $F$ is monotone on $C$ and continuous on $\mathds{R}^{n}$.
Observe that $\partial g_{i}(x)\neq\emptyset$ for $x\in\mathds{R}^{n}$ for all $i\in\{1,\ldots,m\}$, because we assume that $g_{i}$ is convex and $\operatorname{dom}(g_{i})=\mathds{R}^{n}$, for all $i\in\{1,\ldots,m\}$. First, we establish that Algorithm 2 is well-defined.
@@ 4.1. Analysis of the direct method
  2. (ii)
    If $z^{k+1}=\tilde{y}^{k}$, then $\tilde{y}^{k}\in\operatorname{SOL}(F,C)$;
  3. (iii)
    $j(k)$ is well-defined.
@@
  1. (i)
    By (4.1) and Proposition 2.3, we know that $y^{k,j+1}=T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j})\in S_{i}^{k}$ for all $i\in\{1,\ldots,m\}$. In view of the definition of $S_{i}^{k}$ and $g$, we know that (i) holds automatically.
  2. (ii)
    Suppose that $z^{k+1}=\tilde{y}^{k}$. Then, by Proposition 2.3, we get that $\mathbf{\tilde{y}}^{k}\in\mathbf{H}_{k}\subset\mathbf{S}_{k}\subset\mathds{R}^{nm}$. Therefore, $\tilde{y}^{k}\in S_{i}^{k}$ for all $i\in\{1,\ldots,m\}$. Given $i\in\{1,\ldots,m\}$, we have two cases, according to the definition of $S_{i}^{k}$ in (2.4). If $S_{i}^{k}=C_{i}$, then we have $z^{k+1}=\tilde{y}^{k}\in C_{i}$. Otherwise, we have
    $g_{i}(\tilde{y}^{k})=g_{i}(z^{k+1})\leq g_{i}(\tilde{y}^{k})+\langle\tilde{v}_{i}^{k,j},z^{k+1}-\tilde{y}^{k}\rangle\leq 0,$
    where $\tilde{v}_{i}^{k,j}\in\partial g_{i}(\tilde{y}^{k})$. The equality holds by the assumption, the first inequality follows from the subgradient inequality and the last one in consequence of the fact that $\tilde{y}^{k}\in S_{i}^{k}$. Therefore, $\tilde{y}^{k}\in C_{i}$. By the above argument above, we conclude that for any $i\in\{1,\ldots,m\}$, we have $\tilde{y}^{k}\in C$. Moreover, since $z^{k+1}$ is given by (4.5), we can use the similar argument as in Lemma 3.2 to conclude that $\tilde{y}^{k}\in{\rm SOL}(F,C)$.
  3. (iii)
    Assume by contradiction that
    $\frac{g(y^{k,j})\|y^{k,j}-w\|}{g(y^{k,j})-g(w)}>\theta\beta_{k},$
    for all $j$. Thus, we obtain an infinite sequence $\{y^{k,j}\}$ such that
    (4.8) $\liminf_{j\rightarrow\infty}\frac{g(y^{k,j})\|y^{k,j}-w\|}{g(y^{k,j})-g(w)}\geq\theta\beta_{k}>0.$
    Taking into account the inner loop in $j$ given in (4.1) i.e., $y^{k,j+1}\coloneqq T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j})$ for each fixed $k$, we obtain, for each $x\in C$,
    (4.9) $\begin{split}\|y^{k,j+1}-x\|^{2}&=\|T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j})-T_{\mathbf{S}_{k},\mathbf{D}}(x)\|^{2}=\|P_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j})-P_{\mathbf{S}_{k},\mathbf{D}}(x)\|\\ &\leq\|y^{k,j}-x\|^{2}-\|y^{k,j+1}-y^{k,j}\|^{2}\\ &\leq\|y^{k,j}-x\|^{2},\end{split}$
    where the first equality follows from Proposition 2.3 and the first inequality holds because of the firm nonexpansiveness of the projection operator. Thus, $\{y^{k,j}\}_{j\geq 0}$ is quasi-Fejér convergent to $C$, and hence it is bounded by Proposition 2.15. It follows that
    (4.10) $\tau\coloneqq\frac{1}{-g(w)}\sup_{0\leq j\leq\infty}\|y^{k,j}-w\|,$
@@
    (4.12) $\lim_{j\rightarrow\infty}\|y^{k,j+1}-y^{k,j}\|=\lim_{j\rightarrow\infty}\|T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j})-y^{k,j}\|=0.$
@@
    (4.14) $\liminf_{j\rightarrow\infty}g(y^{k,j})\leq 0.$
@@
    (4.15) $\begin{split}\liminf_{j\rightarrow\infty}\frac{g(y^{k,j})\|y^{k,j}-w\|}{g(y^{k,j})-g(w)}&\leq\liminf_{j\rightarrow\infty}\frac{g(y^{k,j})\|y^{k,j}-w\|}{-g(w)}\\ &\leq\frac{1}{-g(w)}\sup_{0\leq j\leq\infty}\|y^{k,j}-w\|\liminf_{j\rightarrow\infty}g(y^{k,j})\\ &=\tau\liminf_{j\rightarrow\infty}g(y^{k,j})\leq 0,\end{split}$
    contradicting (4.8). Therefore, $j(k)$ is well-defined.
@@
If $\operatorname{SOL}(F,C)$ is nonempty, then $\{\tilde{y}^{k}\}$ and $\{z^{k}\}$ are quasi-Fejér convergent to $\operatorname{SOL}(F,C)$.
@@
(4.17) $\frac{\|F(\tilde{y}^{k})\|}{\eta_{k}}\leq 1.$
Take any $x^{*}\in\operatorname{SOL}(F,C)$, we get
(4.18) $\begin{split}\|\tilde{y}^{k}-x^{*}\|^{2}&=\|y^{k,j(k)}-x^{*}\|^{2}=\|T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j(k-1)})-T_{\mathbf{S}_{k},\mathbf{D}}(x^{*})\|^{2}\leq\|y^{k,j(k)-1}-x^{*}\|^{2}\\ &=\|T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j(k)-2})-T_{\mathbf{S}_{k},\mathbf{D}}(x^{*})\|^{2}\leq\|y^{k,j(k)-2}-x^{*}\|^{2}\\ &\cdots\\ &\leq\|y^{k,1}-x^{*}\|^{2}=\|T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,0})-T_{\mathbf{S}_{k},\mathbf{D}}(x^{*})\|^{2}\\ &\leq\|y^{k,0}-x^{*}\|^{2}=\|z^{k}-x^{*}\|^{2}.\end{split}$
@@
(4.19) $\begin{split}\|\tilde{y}^{k+1}-x^{*}\|^{2}&\leq\|z^{k+1}-x^{*}\|^{2}=\left\|T_{\mathbf{S}_{k},\mathbf{D}}(\tilde{y}^{k}-\frac{\beta_{k}}{\eta_{k}}F(\tilde{y}^{k}))-T_{\mathbf{S}_{k},\mathbf{D}}(x^{*})\right\|^{2}\\ &\leq\|\tilde{y}^{k}-\frac{\beta_{k}}{\eta_{k}}F(\tilde{y}^{k})-x^{*}\|^{2}\\ &=\|\tilde{y}^{k}-x^{*}\|^{2}+\frac{\|F(\tilde{y}^{k})\|^{2}}{\eta_{k}^{2}}\beta_{k}^{2}-2\frac{\beta_{k}}{\eta_{k}}\langle F(\tilde{y}^{k}),\tilde{y}^{k}-x^{*}\rangle\\ &\leq\|\tilde{y}^{k}-x^{*}\|^{2}+\beta_{k}^{2}-2\frac{\beta_{k}}{\eta_{k}}\langle F(x^{*}),\tilde{y}^{k}-x^{*}\rangle\\ &=\|\tilde{y}^{k}-x^{*}\|^{2}+\beta_{k}^{2}-2\frac{\beta_{k}}{\eta_{k}}(\langle F(x^{*}),\tilde{y}^{k}-P_{C}(\tilde{y}^{k})\rangle+\langle F(x^{*}),P_{C}(\tilde{y}^{k})-x^{*}\rangle)\\ &\leq\|\tilde{y}^{k}-x^{*}\|^{2}+\beta_{k}^{2}-2\frac{\beta_{k}}{\eta_{k}}\|F(x^{*})\|\|P_{C}(\tilde{y}^{k})-\tilde{y}^{k}\|\\ &\leq|\tilde{y}^{k}-x^{*}\|^{2}+\tilde{\theta}\beta_{k}^{2}\\ &\leq\|z^{k}-x^{*}\|^{2},\end{split}$
using (4.18) in the last inequality, the nonexpansiveness of $T_{\mathbf{S}_{k},\mathbf{D}}$ in the second one, the monotonicity of $F$ and (4.17) in the third one, the fact that $x^{*}\in\operatorname{SOL}(F,C)$ in the fourth one, the Cauchy-Schwartz inequality in the fifth one, Lemma 2.18 and definition of $j(k)$ in the sixth one, and (4.18) in the last one.
In view of Definition 2.12, we conclude that $\{\tilde{y}^{k}\}$ and $\{z^{k}\}$ are quasi-Fejér convergent to $\operatorname{SOL}(F,C)$. ∎
@@
Let $\{z^{k}\}$, $\{\tilde{y}^{k}\}$ and $\{x^{k}\}$ be the sequences generated by Algorithm 2. Assume that $\operatorname{SOL}(F,C)$ is nonempty. Then
  1. (i)
@@
  3. (iii)
    $\lim_{k\rightarrow\infty}\operatorname*{dist}(x^{k},C)=0$;
  4. (iv)
@@
    (4.20) $\begin{split}\sigma_{k}x^{k+1}=\left(\sigma_{k}-\frac{\beta_{k}}{\eta_{k}}\right)x^{k}+\frac{\beta_{k}}{\eta_{k}}\tilde{y}^{k}&=\sigma_{k-1}x^{k}+\frac{\beta_{k}}{\eta_{k}}\tilde{y}^{k}\\ &=\sigma_{k-2}x^{k-1}+\frac{\beta_{k-1}}{\eta_{k-1}}\tilde{y}^{k-1}+\frac{\beta_{k}}{\eta_{k}}\tilde{y}^{k}\\ &\cdots\\ &=\sigma_{0}x^{1}+\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}\tilde{y}^{j}\\ &=\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}\tilde{y}^{j},\end{split}$
@@
  2. (ii)
    For $\{\tilde{y}^{k}\}$ we use Lemma 4.3 and Proposition 2.15(i). For $\{F(\tilde{y}^{k})\}$, we use the boundedness of $\{\tilde{y}^{k}\}$ and the fact that $F$ is monotone and continuous. We use the boundedness of $\{\tilde{y}^{k}\}$ and (i) in for $\{x^{k}\}$.
  3. (iii)
@@
    (4.21) $\operatorname*{dist}(\tilde{y}^{k},C)\leq\theta\beta_{k}.$
@@
    (4.22) $\tilde{x}^{k+1}\coloneqq\frac{1}{\sigma_{k}}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}P_{C}(\tilde{y}^{k}).$
@@
    (4.23) $\tilde{\beta}\coloneqq\sum_{j=0}^{\infty}\beta_{j}^{2}.$
@@
    (4.24) $\begin{split}\operatorname*{dist}(x^{k+1},C)&\leq\|x^{k+1}-\tilde{x}^{k+1}\|=\left\|\frac{1}{\sigma_{k}}\left(\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}(\tilde{y}^{j}-P_{C}(\tilde{y}^{j}))\right)\right\|\\ &\leq\frac{1}{\sigma_{k}}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}\|\tilde{y}^{j}-P_{C}(\tilde{y}^{j})\|\\ &\leq\frac{\theta}{\sigma_{k}}\sum_{j=0}^{k}\frac{\beta_{j}^{2}}{\eta_{j}}\leq\frac{\theta}{\sigma_{k}}\sum_{j=0}^{k}\beta_{j}^{2}\leq\theta\frac{\tilde{\beta}}{\sigma_{k}},\end{split}$
    using the fact that $\tilde{x}^{k+1}\in C$ in the first inequality, (i) and (4.22) in the first equality, convexity of $\|\cdot\|$ in the second inequality, (4.21) in the third one, (4.16) in the fourth one, and (4.23) in the last one.
@@
    (4.25) $\lim_{k\rightarrow\infty}\sigma_{k}=\lim_{k\rightarrow\infty}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}\geq\lim_{k\rightarrow\infty}\frac{1}{\gamma}\sum_{j=0}^{k}\beta_{j}=\infty,$
    using the fact that $\eta_{j}=\max\{1,\|F(\tilde{y})\|\}\leq\max\{1,\gamma\}\leq\gamma$ for all $j$ in the inequality, and (1.22) in the first equality. Thus, taking limits in (4.24), we get
    $\lim_{k\rightarrow\infty}\operatorname*{dist}(x^{k},C)=0,$
    establishing (iii).
  4. (iv)
    Item (iv) follows immediately from (iii).
Lemma 4.5.
If $F$ and $C$ satisfy Assumptions 1.5, 1.8 and 1.9, (i.e. $\operatorname{SOL}(F,C)$$\neq\emptyset$ and $F$ is continuous and monotone), then all the cluster points of $\{x^{k}\}$ generated by Algorithm 2 are solutions of $\operatorname{VIP}(F,C)$.
@@
(4.26) $\begin{split}\|z^{j+1}-x\|&=\left\|T_{\mathbf{S}_{k},\mathbf{D}}\left(\tilde{y}^{j}-\frac{\beta_{j}}{\eta_{j}}F(\tilde{y}^{j})\right)-T_{\mathbf{S}_{k},\mathbf{D}}(x)\right\|\leq\left\|\left(\tilde{y}^{j}-\frac{\beta_{j}}{\eta_{j}}F(\tilde{y}^{j})\right)-x\right\|\\ &=\|\tilde{y}^{j}-x\|^{2}+\frac{\|F(\tilde{y}^{j})\|}{\eta_{j}^{2}}\beta_{j}^{2}-2\frac{\beta_{j}}{\eta_{j}}\langle F(\tilde{y}^{j}),\tilde{y}^{j}-x\rangle\\ &\leq\|\tilde{y}^{j}-x\|^{2}+\beta_{j}^{2}+2\frac{\beta_{j}}{\eta_{j}}\langle F(x),x-\tilde{y}^{j}\rangle,\end{split}$
using nonexpansiveness of $T_{\mathbf{S}_{k},\mathbf{D}}$ in the first inequality, and monotonicity of $F$ and (4.17) in the last inequality. Summing (4.26) from 0 to k1k-1 and dividing by $\sigma_{k-1}$, we obtain, in view of Lemma 4.4(ii),
(4.27) $\frac{(\|z^{k}-x\|^{2}-\|z^{0}-x\|^{2})}{\sigma_{k-1}}\leq\frac{\sum_{j=0}^{k-1}\beta_{j}^{2}}{\sigma_{k-1}}+\langle F(x),x-x^{k}\rangle.$
@@
for all $x\in C$. Using Proposition 2.14, we get that $\hat{x}\in\operatorname{SOL}(F,C)$. Therefore, all the cluster points of $\{x^{k}\}$ solve $\operatorname{VIP}(F,C)$. ∎
We complete next the convergence analysis of the algorithm.